7(d-5)+1=5 What is d??
Answer by checkley71(8403)

7(D-5)+1=5
7D-35+1=5
7D=5+35-1
7D=39
D=39/7
PROOF
7(39/7-5)+1=5
7(39/7-35/7)+1=5
7(4/7)+1=5
28/7+1=5
4+1=5
5=5